How can i find the x-intercept and y-intercept of the line given by the equation 5x-7y=15
Answer by lwsshak3(11628)

How can i find the x-intercept and y-intercept of the line given by the equation 5x-7y=15
**
To find the x-intercept, set y=0, then solve for x.
To find the y-intercept, set x=0, then solve for y
..
5x-7y=15
set y=0
5x=15
x=3 (x-intercept)
..
5x-7y=15
set x=0
-7y=15
y=-15/7 (y-intercept)